Question 1 (6 points) Two college instructors are interested in whether or not there is any variation in the way they grade mQuestion 2 (3 points) Based on our p-value, compared with a 5% alpha level we should Oreject the null hypothesis that all popBasketball ANOVA Four basketball teams took a sample of 3 best players regarding how high each player can jump (in inches). TQuestion 8 (3 points) Based on our p-value, compared with a 5% alpha level we should reject the null hypothesis that all popuPlease answer all 9 questions
Answer #1

Page No. Date Sg.2 = 72.5. :. Ftest test shahrtic = NI = 12.5 59.2 - 1.225 [1.225 Here P value = o.osgaza o-059229 > 0.05. T:Page No. Date 0.3 F critical = 3.885294 Ftest (calculated) = 3.610133 .. Here, Ecal feritical. - Decision Rule! Ang! fail toDate 1.51 least variation Group A. Which team hus least varighion Team 5 Team s - a.7 closest menn Team I-Team 4 = 43- 42.666Date Q. g) Ecritical value = 3.47805 F calculated value = 2.060393. . E calculated value < Fcritical value L. fail to reject